OpenStudy (anonymous):

what is the discriminant of x2 - 14x + 49 = 0

OpenStudy (anonymous):

I am not sure if it is 392 or 0. do I add or subtract??

OpenStudy (anonymous):

14^-4(49)=0... you subtract

OpenStudy (anonymous):

really? ok, well thank you. so there is really only one solution, correct?

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

thanks so much, that what i thought. :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

disc= b^2 - 4ac here a= 1 , b= -14 and c= 49 so disc = (-14)^2 - 4(1)(49) disc= 196- 196 disc = 0